# Initial Guessing Bias: How Untrained Networks Favor Some Classes

### v4

@@ I.4 Reproducibility
        a first convolutional layer with: out channels=16 (Number of channels produced by the convolution), $m=5$ ( Size of the convolving kernel), stride=1, padding=2. The output of the layer is then passed through an activation function and a pooling layer. The choice of these elements varies to compare two different scenarios (as explained in Fig. 1). In both cases, however, we set common parameters for the pooling layer, i.e. $m=2$ (kernel size), stride=2.
        Next comes a second convolutional layer with the same parameters as the previous one, except for the number of output chanels; in this case we have out channels=64. Again the convolutional layer is followed by an activation function and a pooling layer (same as the first layer). The parameters of the pooling layer in this case are i.e. $m=4$ (kernel size), stride=4. The processed signal is then connected with a weights layer to the output layer.
@@
        we start with a sequence of five convolutional layers (each followed by an activation function and pooling layer). Except for the number of output channels the rest of the parameters are fixed the same for each of these layers, in particular we have for the convolutional layer =5=5, stride=1, padding=4. For the pooling layer, however, $m=5$, stride=1. Finally, the number of output channels for the various layers is $[16,32,32,64,32]$.
        This is followed by an additional convolutional layer defined by the following parameters: out channels=16, $m=5$, stride=1, padding=2. This is accompanied by the activation function and a Pooling layer whose parameters are: $m=2$, stride=2.
        Finally, a last sequence of convolutional layers, activation function and pooling layer precedes the output layer. The only parameter that differs from the sequence that precedes it is the kernel size of the pooling layer; specifically in this case $m=4$. The processed signal is then connected with a weights layer to the output layer.

### v5

@@ I.4 Reproducibility
        a first convolutional layer with: out channels=16 (Number of channels produced by the convolution), $k=5$ ( Size of the convolving kernel), stride=1, padding=2. The output of the layer is then passed through an activation function and a pooling layer. The choice of these elements varies to compare two different scenarios (as explained in Fig. 1). In both cases, however, we set common parameters for the pooling layer, i.e. $k=2$ (kernel size), stride=2.
        Next comes a second convolutional layer with the same parameters as the previous one, except for the number of output chanels; in this case we have out channels=64. Again the convolutional layer is followed by an activation function and a pooling layer (same as the first layer). The parameters of the pooling layer in this case are i.e. $k=4$ (kernel size), stride=4. The processed signal is then connected with a weights layer to the output layer.
@@
        we start with a sequence of five convolutional layers (each followed by an activation function and pooling layer). Except for the number of output channels the rest of the parameters are fixed the same for each of these layers, in particular we have for the convolutional layer =5=5, stride=1, padding=4. For the pooling layer, however, $k=5$, stride=1. Finally, the number of output channels for the various layers is $[16,32,32,64,32]$.
        This is followed by an additional convolutional layer defined by the following parameters: out channels=16, $k=5$, stride=1, padding=2. This is accompanied by the activation function and a Pooling layer whose parameters are: $k=2$, stride=2.
        Finally, a last sequence of convolutional layers, activation function and pooling layer precedes the output layer. The only parameter that differs from the sequence that precedes it is the kernel size of the pooling layer; specifically in this case $k=4$. The processed signal is then connected with a weights layer to the output layer.